Shift Summary for DAY, Aug.21, 1999

Angela Bellavance (bellavan@rice.edu)
Sat, 21 Aug 1999 15:41:45 -0500 (CDT)

Shift Summary for day shift, 21/Aug/99

On Shift: D.Jensen, A.Bellavance

Hours of beam delivered at all intensities: 8.4 (of 8.5)
Hours of beam delivered at nominal intensity: 8.4
Hours of E832 data (special runs not included): 7.4
Hours of special runs (laser scans, muon runs...): 0

AMBNote: OK, so I put too many significant digits in yesterday. I was
translating from minutes to decimal as requested by the run manager. No
need for rude comments please. Sheesh...

Accounting of lost time: changing tapes, pipeline problems

========================================================================
Runs ended this shift:
Run 14291 - 23.7M events - BEAM_832
Run 14292 - ???? events - BEAM_832 (in progress)

Tapes brought to FCC this shift: QKC297-299,
QKK297-299,
QKS297-299
QKB104

========================================================================
Accesses during shift:none
========================================================================
Overview of shift: E832 datataking, a few pipeline problems
========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
14291 311 7:00 Take over w/beam
7:05 Tape mounted: QKNL24
355 Ktevana errors - KFREND
~7:45 Notice index 226 has gain drift -5%. Ed wants to
be paged when tapes fill up to update constants.
356 8:03 Beam abort
368 8:18 Trigger dies: DYC full in laser crate - KFREND
14291 385 8:41 Tapes fill up. Ed paged and he updates constants.
14292 1 9:01 Start run
9:05 Tape mounted: QKCR06
18 Ktevana errors - KFREND
27 Start getting occasional messages on the Serious
Alarms screen: "Message from pipeline", then
nothing else. Decide to just click "ok".
46 Beam disappears for one spill.
140-1 12:08 Ktevana errors - KFREND
160 Laser RMS Stab is 29% on channel 2219. Corrects
itself by next spill.
12:45 RickK calls about RC10. trg#veto looks bad but
rcmon#srctime is happy. Keep an eye on it, may
need to be fixed.
177 12:57 Ped. drift plot shows a vertical ribbon in the
lower right corner. Julie notified and is looking
at it.
185-207 Problem looks like a pipeline error. Stop
triggers to re-init with DBCs. Crate 7 is not
responding. Reboot crate 7. Init hangs on crate
10. Reboot crates 10 and 11. Init works. Did 2
KFRENDS. Restarted triggers but problem still
shows up. Suspect the problem was in the first
spill of the run and so won't go away. Leave it
alone for now but check at beginning of next run.
The "Message from pipeline" error window
does not appear again after this.
214 Ktevana errors - KFREND.
14:05 cosmic#summary plots (Y-track finder inputs) have
spikes on them. RickK notified by email.
14:12 fpedcal plots show a few channels drifting.
Valeri notified by email.
14:53 Tape mounted: QKSM13
269 Laser stab. on #2371=34% and something about
"DAQ_SEC" being dead. Ran a KFREND, seems to have
worked.
270-272 Laser stab. on #558=5%. Settles back down on its
own.
273 15:06 Beam abort for one spill.
289 Beam gone for one spill.
290-? Beam starts aborting at the end of each spill.
Hand off to eve shift.

========================================================================
Follow up Notes on Previous Shift: none
========================================================================
Items to Follow up on Next Shift:

- CsI channel 226 gain drift has been slowly drifting down. Keep an eye
on it.

- RC10 is having some trouble (bad trg#veto plot) but as long as the
rcmon#srctime plot is happy, we are ok. Keep an eye on it. (RickK)

-check pipeline crates (cped#drlist) at beginning of next run.

_______
Angela Bellavance work: (630) 840-4144
bellavan@rice.edu home: (630) 548-3412
http://www.owlnet.rice.edu/~bellavan/ page: (630) 722-6861
"It has yet to be proven that intelligence has any survival value."
- Arthur C. Clarke